OpenStudy (anonymous):

if 0 = (16x-x) does it mean x=1/16?

OpenStudy (dumbcow):

no 16x-x = 15x ...x = 0

OpenStudy (dumbcow):

0 = 16x - 1 --> x = 1/16

OpenStudy (lgbasallote):

the dumbcow is right :)

OpenStudy (anonymous):

\[\log _{2}x-\log _{4}x = -2 \] how then? :/

OpenStudy (lgbasallote):

tricky...you need to use change-of-base formula..are you familiar with that?